# Gradient Reweighting: Towards Imbalanced Class-Incremental Learning

### v1

@@ 3.1 Imbalanced Gradients
Imbalanced gradient updates pose a significant challenge when learning from imbalanced data. Generally, consider the weights of a classifier with input feature dimension and output classes . Given a loss function , the gradient-based weight update process is formulated as
@@ 4.1 Intra-Phase Class-Imbalance
Regularized softmax cross-entropy. While the gradient update adjustments help in emphasizing tail classes, it also inadvertently leads to increased gradients towards head classes due to the decrease of output logit of head classes, making the cross-entropy loss put more efforts to increase its output logits. Motivated by [33], we alleviate this side effect by applying regularized cross-entropy using a modified softmax equation
where is the per-label offsets added to the original output logit. Generally, the instance-rich classes have larger with an increase of softmax output to compensate for the side effect of down-weighting the gradients during the training process. In this work, we consistently use the accumulated gradient to calculate at each iteration , which ensures a balanced training by adjusting the output logit for head classes to mitigate the over-amplified efforts.
@@ 4.2.2 Decoupled Gradient Reweighting
For tuning between stability and plasticity, we adjust the magnitude of gradients from distillation to make it balance with reweighted cross-entropy by including a loss balance ratio at iteration as . The overall DGR can be formulated as
@@ 5.1 Experimental Setups
Evaluation protocol. In this work, we adopt two widely used protocols for CIL including (1) learning from scratch (LFS) [37] and (2) learning from half (LFH) [20, 14]. The LFS equally split all the classes into tasks to incrementally train a model from scratch. The LFH first trains the model with the initial half of the classes and then equally divides the remaining half of the classes into tasks. For both cases, the model is evaluated on all classes seen so far after learning each task without knowing the task identifier.
@@
Implementation details. We utilize ResNet-32 [18] for CIFAR100-LT and ResNet-18 [18] for ImageNetSubset-LT and Food101-LT. To ensure a fair comparison, we adopt the same class order and setups from [26] where we train 160 epochs for CIFAR100-LT with an initial learning rate of 0.1 and then reduced by a factor of 10 at the 80th and 120th epochs. For ImageNetSubset-LT and Food101-LT, we train 90 epochs with an initial learning rate of 0.1 which is reduced by 10 at the 30th and 60th epoch. A consistent batch size of 128 with an SGD optimizer is used across all experiments. For simplicity, we set our hyper-parameters with and for all experiments. We store exemplars per class selected by Herding [37] and conduct each experiment three times to report average performance. All results are obtained by reproducing the original methods under the same setting based on the framework in [31, 26].
@@ 5.2 Experimental Results
In Figure 4, we visualize the classification accuracy on all classes seen so far at each incremental step by varying the imbalance factor and the memory budget . While generally increasing the memory budget results in better performance, the improvements could be marginal for some methods such as SSIL [1] and MAFDRC [11]. This paradox is attributed to the exacerbation of exemplar set imbalance by using a larger memory budget where the instance-rich classes retain more exemplars than instance-rare classes. On the other hand, the increase of imbalance factor results in a noticeable degradation in performance. Despite these challenges, our method is able to manage diverse learning environments posed by both memory budget and imbalance factors to consistently achieve the best performance at each incremental step. Additional results can be found in supplementary materials.
@@ 5.3 Ablation Study
In this section, we evaluate the effectiveness of our method to address (i) intra-phase imbalance, (ii) inter-phase imbalance, and (iii) biased weights in FC layers.
Intra-phase Imbalance: We evaluate our gredient reweighting for its efficacy in solving the imbalanced image classification beyond CIL by conducting comparative analyses with established methods in long-tailed recognition. We denote our regularized softmax output illustrated in Section 4.1 as RS. Specifically, we train a ResNet-32 to classify the 100 classes in CIFAR100-LT with various imbalance factor following the training protocol as in [13]. The results are summarized in Table 2. Our method shows competitive performance even without the use of regularized softmax output. Upon integrating the regularized softmax output, which further balances the learning process, we consistently achieve improved classification accuracy and outperform existing methods.
CIFAR100-LT | |||
ROS [43] | 36.32 | 41.28 | 55.12 |
Focal Loss [25] | 38.91 | 43.26 | 55.08 |
LDAM [9] | 40.82 | 45.68 | 57.32 |
CB Loss [13] | 39.62 | 46.29 | 57.29 |
IB Loss [34] | 43.62 | 46.80 | 58.01 |
BS Loss [38] | 44.12 | 49.25 | 59.38 |
EQL v2 [42] | 43.81 | 48.25 | 57.06 |
CMO [35] | 43.54 | 47.92 | 58.97 |
Ours w/o RS | 43.84 | 47.39 | 57.95 |
Ours | 45.27 | 49.02 | 60.71 |
Inter-phase Imbalance: We evaluate each of our components including (a) the addition of our DAKD loss, (b) the Decomposed Gradient Reweighting (DGR) to separately balance the gradients for new and learned tasks instead of treating all classes together (GR) as for intra-phase imbalance, and (3) the significance of maintaining the discrimination of gradients from knowledge distillation loss ((DAKD)) as described in Section 4.2.2. The results are summarized in Table 3, where we observe performance improvements contributed by each component. Notably, the GR exhibits inferior performance compared to DGR due to the potential underfitting of new classes which received insufficient emphasis during CIL. We also observe improved performance by maintaining the discriminativeness of the gradient from DAKD to capture and leverage information from the original training distribution, thus effectively preserving learned knowledge.
Additionally, we assess the effectiveness of DAKD to mitigate forgetting under CIL by replacing it with existing variants of logit-based knowledge distillation loss including (i) the balanced knowledge distillation (BKD) [51], (ii) multi-level logit distillation(MKD) [21], (iii) decoupled knowledge distillation (DCKD) [54], and (iv) decomposed knowledge distillation(DPKD) [4]. The results are visualized in Figure 5 () where we measure the forgetting rate [29] at each incremental learning phase. Remarkably, we notice that the MKD, DCKD, and BKD achieve comparable or even worse results compared to the original KD loss as they assume the student and teacher model observe the same training distribution, which does not hold under CIL. While DPKD marks some improvements, its effectiveness remains hampered by imbalanced data. Overall, our DAKD achieves the lowest forgetting and best ACC by factoring lost training data distribution.

### v2

@@ 3.1 Imbalanced Gradients
Imbalanced gradient updates pose a significant challenge when learning from imbalanced data as studied in [15]. Generally, consider the weights of a classifier with input feature dimension and output classes . Given a loss function , the gradient-based weight update process is formulated as
@@ 4.1 Intra-Phase Class-Imbalance
Regularized softmax cross-entropy. While the gradient update adjustments help in emphasizing tail classes, it also inadvertently leads to increased gradients towards head classes due to the decrease of output logit of head classes, making the cross-entropy loss put more effort into increasing its output logits. Motivated by [33], we alleviate this side effect by applying regularized cross-entropy using a modified softmax equation
where denotes the estimates of class priors and is the per-label offsets added to the original output logit. In practice, the calibrated class distribution with the number of data for each seen class is widely used to determine . Generally, the instance-rich classes have larger with an increase of softmax output to compensate for the side effect of down-weighting the gradients during the training process.
@@ 4.2.2 Decoupled Gradient Reweighting
For tuning between stability and plasticity, we adjust the magnitude of gradients from distillation to make it balance with reweighted cross-entropy by including a loss balance ratio at iteration as
The overall DGR can be formulated as
@@ 5.1 Experimental Setups
Evaluation protocol. In this work, we adopt two widely used protocols for CIL including (1) learning from scratch (LFS) [37] and (2) learning from half (LFH) [21, 14]. The LFS equally split all the classes into tasks to incrementally train a model from scratch. The LFH first trains the model with the initial half of the classes and then equally divides the remaining half of the classes into tasks. For both cases, the model is evaluated on all classes seen so far after learning each task without knowing the task identifier. We apply growing memory with exemplars per class selected by Herding [37] for both LFS and LFH. The discussion about fixed memory setup can be found in Appendix.
@@
Implementation details. We utilize ResNet-32 [19] for CIFAR100-LT and ResNet-18 [19] for ImageNetSubset-LT and Food101-LT. Both feature extractor and classifiers are trained from scratch using Equation 9 in an end-to-end fashion. To ensure a fair comparison, we adopt the same class order and setups from [26] where we train 160 epochs for CIFAR100-LT with an initial learning rate of 0.1 and then reduced by a factor of 10 at the 80th and 120th epochs. For ImageNetSubset-LT and Food101-LT, we train 90 epochs with an initial learning rate of 0.1 which is reduced by 10 at the 30th and 60th epoch. A consistent batch size of 128 with an SGD optimizer is used across all experiments. For simplicity, we set our hyper-parameters with and for all experiments. Each experiments are conducted three times to report average performance. All results are obtained by reproducing the original methods under the same setting based on the framework in [31, 26].
@@ 5.2 Experimental Results
In Figure 4, we visualize the classification accuracy on all classes seen so far at each incremental step by varying the imbalance factor and the memory budget . While generally increasing the memory budget results in better performance, the improvements could be marginal for some methods such as SSIL [1] and MAFDRC [11]. This paradox is attributed to the exacerbation of exemplar set imbalance by using a larger memory budget where the instance-rich classes retain more exemplars than instance-rare classes. On the other hand, the increase of imbalance factor results in a noticeable degradation in performance. Despite these challenges, our method is able to manage diverse learning environments posed by both memory budget and imbalance factors to consistently achieve the best performance at each incremental step.
Additional results for (i) conventional CIL setup (), (ii) long-tail recognition, and (iii) computation and memory evaluations can be found in Appendix.
@@ 5.3 Ablation Study
We evaluate each of our components including (i) the addition of our DAKD loss, (ii) the Decomposed Gradient Reweighting (DGR) to separately balance the gradients for new and learned tasks instead of treating all classes together (GR) for intra-phase imbalance, and (iii) the significance of maintaining the discrimination of gradients from knowledge distillation loss ((DAKD)) as described in Section 4.2.2. Specifically, we consider GR and original knowledge distillation (KD) as the baseline to evaluate the contributions by adding each component from our method. The results are summarized in Table 2, where we observe performance improvements contributed by each component. Notably, the GR exhibits better performance compared to DGR due to the potential underfitting of new classes which received insufficient emphasis during CIL. We also observe improved performance by maintaining the discriminativeness of the gradient from DAKD to capture and leverage information from the original training distribution, thus effectively preserving learned knowledge.
Variants of knowledge distillation loss: Additionally, we assess the effectiveness of DAKD to mitigate forgetting under CIL by replacing it with existing variants of logit-based knowledge distillation loss including (i) the balanced knowledge distillation (BKD) [51], (ii) multi-level logit distillation(MKD) [22], (iii) decoupled knowledge distillation (DCKD) [54], and (iv) decomposed knowledge distillation(DPKD) [4]. The results are visualized in Figure 5 () where we measure the forgetting rate [29] at each incremental learning phase. Remarkably, we notice that the MKD, DCKD, and BKD achieve comparable or even worse results compared to the original KD loss as they assume the student and teacher model observe the same training distribution, which does not hold under CIL. While DPKD marks some improvements, its effectiveness remains hampered by imbalanced data. Overall, our DAKD achieves the lowest forgetting and best ACC by factoring lost training data distribution.
@@ 6 Conclusion
Acknowledgement
This project was supported by the National Institutes of Health (Grant No. U24CA268228). Special thanks are extended to Professor Fengqing Zhu for invaluable comments and advice on the manuscript, with a regrettable oversight of not being included in the author list.